OpenStudy (anonymous):

What is the length of the mid segment of the trapezoid made by the vertices A(0,5) B(3,3) C(5,-2) and D(-1,2). Show equations and all work that leads to your answer. Help Please:s

OpenStudy (anonymous):

@SolomonZelman Helpp Please

OpenStudy (solomonzelman):

That's a lot of work...

OpenStudy (anonymous):

Indeed it is :s

OpenStudy (solomonzelman):

Distance formula. \[d=\sqrt{(y_1-y_2)^2+(x_1-x_2)^2}\] I would suggest drawing it first.

OpenStudy (solomonzelman):

(If I make an attachment, it would take a while.)

OpenStudy (solomonzelman):

I don't have a right program to draw it, can you draw it?

OpenStudy (anonymous):

distance formula d² = ( x2- x1)² + ( y2-y1)² segment AB D²(AB)= (3-0)² + (3-5)²= 9+4=13 D AB = √ 13 midpoint will divide the segment in half, thus each half has a length of D (AB)/ 2 = √(13)/2 I dont know

OpenStudy (anonymous):

I will try and see if If theres one on the internet, wait a sec